### v1

# Memory Gym: Partially Observable Challenges to Memory-Based Agents in Endless Episodes

Abstract
Memory Gym introduces a unique benchmark designed to test Deep Reinforcement Learning agents, specifically comparing Gated Recurrent Unit (GRU) against Transformer-XL (TrXL), on their ability to memorize long sequences, withstand noise, and generalize. It features partially observable 2D environments with discrete controls, namely Mortar Mayhem, Mystery Path, and Searing Spotlights. These originally finite environments are extrapolated to novel endless tasks that act as an automatic curriculum, drawing inspiration from the car game “I packed my bag”. These endless tasks are not only beneficial for evaluating efficiency but also intriguingly valuable for assessing the effectiveness of approaches in memory-based agents. Given the scarcity of publicly available memory baselines, we contribute an implementation driven by TrXL and Proximal Policy Optimization. This implementation leverages TrXL as episodic memory using a sliding window approach. In our experiments on the finite environments, TrXL demonstrates superior sample efficiency in Mystery Path and outperforms in Mortar Mayhem. However, GRU is more efficient on Searing Spotlights. Most notably, in all endless tasks, GRU makes a remarkable resurgence, consistently outperforming TrXL by significant margins.
Website and Source Code: https://github.com/MarcoMeter/endless-memory-gym/
Keywords: deep reinforcement learning, proximal policy optimization, memory, transformer-xl, gated recurrent unit
1 Introduction
In the evolving landscape of Deep Reinforcement Learning (DRL), the significance of memory-based tasks is increasingly evident. Our research introduces “Memory Gym”, a benchmark that emphasizes assessing memory capabilities in endless tasks, drawing inspiration from the classic car game “I packed my bag”. Central to our contribution is a Transformer-XL (Vaswani et al., 2017) driven baseline implementation for Proximal Policy Optimization (PPO) (Schulman et al., 2017), making it readily accessible to the community. This introduction delves into the intricate role of memory in decision-making agents, setting the stage for our key contributions. We conclude this section by outlining the structure and flow of the subsequent content in this paper.
1.1 “I Packed My Bag”: A Game to Explore Memory Limits
Imagine embarking on a long-awaited family vacation, with the open road stretching ahead. As the car hums along, a lively atmosphere envelops the passengers, young and old alike, as they engage in a playful game of “I packed my bag”. However, as the game unfolds and the list lengthens, the passengers encounter mounting challenges illustrating the finite nature of their individual memories. Despite best efforts, varying strategies, and different capabilities, the weight of the growing list eventually leads to resignation. Due to the game’s ever-growing or endless concept, it serves as a revealing test of memory effectiveness, showcasing the limitations that individuals face when confronted with the increasing demands of retention and recall.
1.2 Memory in the Context of Decision-Making Agents
Memory is not just a game – it is a critical tool for intelligent decision-making under imperfect information and uncertainty. Without the ability to recall past experiences, reasoning, creativity, planning, and learning may become elusive. In the realm of autonomously learning decision-making agents, the agent’s memory involves maintaining a representation of previous observations, a knowledge bank that grounds its next decision. Memory mechanisms, be it through recurrent neural networks (Rumelhart et al., 1986) or transformers (Vaswani et al., 2017), have enabled these agents to master tasks both virtual and real. For instance, DRL methods have conquered complex video games as Capture the flag (Jaderberg et al., 2018), StarCraft \@slowromancapii@ (Vinyals et al., 2019), and DotA 2 (Berner et al., 2019). Their success extends beyond virtual environments to real-world challenges as dexterous in-hand manipulation (Andrychowicz et al., 2020) and controlling tokamak plasmas (Degrave et al., 2022). The remarkable achievements of DRL agents empowered by memory often come coupled with substantial computational demands and additional techniques that extend beyond the scope of assessing an agent’s memory interaction abilities.
1.3 Contribution: Endless Environment Memory Benchmark
In our prior research, we introduced Memory Gym (Pleines et al., 2023), an open-source benchmark, designed to challenge memory-based DRL agents to memorize events across long sequences, generalize, be robust to noise, and be sample efficient. Memory Gym encompasses three unique environments: Mortar Mayhem, Mystery Path, and Searing Spotlights. Each environment presents visual observations, discrete action spaces, and crucially, they cannot be mastered without memory, demanding frequent memory interactions.
Building on this foundation, we further enhance Memory Gym’s offerings by introducing endless tasks, drawing inspiration from the “I packed my bag” game. Such endless tasks, behaving as an automatic curriculum, are pivotal for not just evaluating efficiency but, more critically, assessing the effectiveness of memory mechanisms. As mirrored in the aforementioned game, capitulating to the ever-expanding list becomes inevitable. To our current understanding, no existing Deep Learning memory benchmark possesses such endless tasks targeting sequence models. Additionally, DRL tasks exhibit dynamic behaviors due to the evolving nature of the agent’s policy during training. As a result, episodes can yield vastly different outcomes. For example, episode lengths can vary significantly based on the decisions made, particularly when dealing with stochastic agents or environments.
It is important to clarify that “endless” should not be confused with “open-ended”. Memory Gym’s endless environments are not open-ended in nature; they possess clear objectives and defined structures. In contrast, open-ended environments offer agents freedom to develop unique skills and strategies without stringent constraints as demonstrated by the Paired Open-Ended Trailblazer algorithm (Wang et al., 2020) or platforms like Minecraft (Fan et al., 2022).
1.4 Contribution: Transformer-XL PPO Baseline
Many significant publications in memory-based DRL, particularly those utilizing transformers (Fortunato et al., 2019; Parisotto et al., 2020; Hill et al., 2021; Lampinen et al., 2021; Parisotto and Salakhutdinov, 2021), withhold their source code, impeding the replication of their findings and consequently hampering research progress. The absence of available baseline implementations is further discussed in section 5.2. In response to this challenge, we offer an accessible and easy-to-follow baseline implementation based on the PPO DRL algorithm (Schulman et al., 2017) and Transformer-XL (TrXL) (Vaswani et al., 2017). To allow for memory, we utilize TrXL as an episodic memory leveraging a sliding window approach.
In our experimental analysis, we benchmark both the TrXL and a recurrent agent based on the Gated Recurrent Unit (GRU) (Cho et al., 2014) across the original (finite) and the new endless environments of Memory Gym. Our findings indicate that TrXL performs best in terms of sample efficiency on Mystery Path and effectiveness on Mortar Mayhem. However, in the Searing Spotlights environment, GRU demonstrates better sample efficiency. To our greatest surprise, in the endless environments, GRU consistently surpasses TrXL in effectiveness by large margins.
1.5 Overview
This paper proceeds by delving into the properties and dynamics of the Memory Gym environments, highlighting Mortar Mayhem, Mystery Path, and Searing Spotlights. Additionally, we introduce their innovative endless counterparts, offering a fresh perspective on memory challenges. Subsequently, we outline the memory baselines that are tested against Memory Gym. Delving deeper, we describe the actor-critic model architecture, shedding light on the associated loss functions and the foundational components of both the GRU and TrXL baselines. Our experimental analysis follows, presenting empirical results from both finite and endless tasks within Memory Gym. Moreover, we discuss the intriguing aspect of TrXL’s suboptimal performance in endless tasks, offering insights, further experimentation, and future directions. Prior to our concluding remarks, we offer a comparative perspective by discussing related baselines and benchmarks in the memory-based DRL domain.
2 Memory Gym Environments
(a) Multi-Discrete
(b) Move Forward
(c) Rotate Left
(d) Rotate Right
Figure 1: The skin-colored agent in Memory Gym utilizes a multi-discrete action space (a) with distinct dimensions for horizontal (red arrows) and vertical (green arrows) velocity. Both dimensions include a no-op action, enabling movement in eight distinct cardinal directions (all arrows). The agent maintains a fixed speed of 3 pixels per step, with its forward direction determined by the combined horizontal and vertical velocity. The agent’s forward direction is visually represented by white-colored hands. Additionally, Mortar Mayhem and Mystery Path offer a grid-like locomotion system with a discrete action space of four actions: moving forward one tile at a time (b), rotating by 90 degrees left (c) or right (d), and a no-op action.
Mini-games of the commercial video game Pummel Party [^1] http://rebuiltgames.com/ inspired Memory Gym’s environments. The agent perceives all environments from a top-down perspective using $84\times 84$ RGB pixels, while its action space is multi-discrete, featuring two dimensions of size three as shown in Figure 1(a). Simpler grid-based locomotion variants are also incorporated in Mortar Mayhem and Mystery Path, as depicted in Figures 1(b) to 1(d). Subsequently, we dive into an in-depth exploration of the environments’ dynamics, highlighting their peculiarities towards memory, and above all, showcasing their distinctive endless design. Annex A presents the reset parameters of the enironments and their default values.
2.1 Mortar Mayhem
[…]
(a) Annotated Ground Truth
(b) Agent Observation
(c) Visual Feedback
Figure 2: Mortar Mayhem’s relevant entities are depicted in the annotated ground truth of the environment (a). The ground truth includes a green circle that designates the next target tile to move to. At the beginning of an episode, the commands are sequentially rendered onto the agent’s observation (b) while the agent remains immobile. Once all commands are displayed, the agent must navigate to the target tile within a specified time frame. Upon completion of the current command’s execution time, the agent’s success is evaluated, as illustrated in figure (c). This visual feedback is perceptible to the agent. Standing on red colored tiles means failure. Following a brief delay of a few steps during evaluation, the agent can proceed to the next command unless the episode has terminated due to failure or completion of the entire command sequence.
Mortar Mayhem (MM; Figure 2) takes place inside a grid-like arena and consists of two memory-dependent tasks. Initially, the agent is immobile and must memorize a sequence of ten commands (Clue Task), followed by executing each command in the observed order (Act Task). A command instructs the agent to move to an adjacent floor tile or remain at the current one. Failure to execute a command results in episode termination, whereas each successful execution yields a reward of $+0.1$. MM can be simplified to solely provide the Act Task (MMAct), where the command sequence is fully observable as a one-hot encoded feature vector. The Act Task necessitates frequent memory utilization since it must discern which commands have been executed to fulfill the subsequent ones, while considering a maximum of nine target tiles. To overcome this challenge, the agent can learn to track time (e.g. count steps) where a short-term memory seems sufficient.
MM’s difficulty can be scaled by several parameters, such as the number of commands, the show duration of a single command, and the delay between showing commands. Furthermore, the time allocated for executing a single command can be modified. These options can also be sampled upon resetting an episode. To simplify the difficulty further, the agent can be equipped with grid-like movement (MMGrid, MMActGrid), allowing it to move one tile at a time. Doing so, the agent is not commanded to move to diagonal tiles.
Episode lengths in MM vary based on the agent’s proficiency, lasting longer as the agent improves until reaching an upper bound (i.e. max episode length). The following equations determine the maximum episode length in MM by considering the durations and delays of the Clue Task and Act Task.
Clue Task $=(\text{Show Duration}+\text{Show Delay})\times\text{Command Count}$ (1)
Execution Length $=(\text{Execution Duration}+\text{Execution Delay})\times\text{Command Count}$ (2)
Act Task $=\text{Execution Length}-\text{Execution Delay}+1$ (3)
[…]
2.2 Endless Mortar Mayhem
(a) Observing and Executing Commands
(b) Agent Observation
Figure 3: In Endless Mortar Mayhem, the visualization and execution of commands are alternated (a) following the idea of an automatic curriculum. Commands (blue) are visualized only once in-between executing a new command and the previous ones. The arena (b) seamlessly occupies the entire screen, enabling the agent to reenter from the opposite side if it walks off.
Endless Mortar Mayhem (EMM; Figure 3) introduces a unique concept inspired by the toddler’s game ”I packed my bag…”. It extends the to-be-executed command sequence continuously, enabling potentially infinite sequences. Two significant modifications are made to MM to accommodate this feature. In the previous setup, the agent had to observe the entire sequence before executing it. However, in EMM, we adopt a different approach by alternating between command visualization and execution.
As shown in Figure 3(a), the agent only perceives one command during visualization. During execution, the agent must execute all previous commands along with the new one. An episode begins by displaying one command, followed by its execution. Subsequently, the next visualization reveals only the second command, requiring the agent to execute both the first and second commands. This automatic curriculum allows the to-be-executed command sequence to continually expand, while each new command is presented only once. A reward of $+0.1$ is signaled for every single successful command.
The second modification concerns the arena’s boundaries. In the original MM, the agent’s position would need to be reset after command execution due to certain commands being unavailable when the agent is adjacent to the arena’s boundaries. In EMM, we eliminate the borders and seamlessly scale the arena to the screen’s dimensions. Additionally, we implement a screen wrap feature that transforms the arena into a torus. Consequently, if the agent walks off the screen, it reappears on the opposite side without requiring relocation. Finally, the agent then executes the sequence of commands again from its new position.
2.3 Mystery Path
[…]
(a) Annotated Ground Truth
(b) Agent Observation
(c) Visual Feedback
Figure 4: The annotated ground truth depicts the relevant entities in Mystery Path (a). The walkable path is distinguished by its blue (origin), green (goal), and white color (path). Areas outside the path are designated as pits, represented by black and red colors, causing the agent to fall off and be relocated to the path’s origin on the next step. The initial randomly chosen red tiles ($A^{\ast}$ Walls) are only considered during path generation. The agent’s observation includes just itself (b). If the agent deviates from the path, a red cross provides visual feedback (c).
Mystery Path (MP; Figure 4) challenges the agent to traverse an invisible path within a grid-like level of dimension $7\times 7$. If the agent deviates from the path and thus falls down the pit, the agent is relocated to its origin during the next step. The episode terminates when the agent reaches the goal or exhausts its time budget of 512 steps. Reaching the goal yields a reward of $+1$. If a dense reward function is desired, the agent can be rewarded by $0.1$ for visiting previously unexplored tiles. To effectively navigate the environment, the agent must remember its steps on the path as well as the locations where it fell off. The invisible path is procedurally generated using the $A^{\ast}$ path-finding algorithm (Hart et al., 1968). Initially, the path’s origin is randomly sampled from the grid’s border, while the goal is placed on the opposing side in a random position. Instead of aiming for the shortest path, the cost function of $A^{\ast}$ is sampled to generate a diverse range of paths. The path generation process randomizes some cells as untraceable ($A^{\ast}$ walls), which adds variation to the generated paths. It is important to note that these walls are considered pits rather than physical obstacles.
MP can be simplified into a grid variant called MPGrid, which features a single discrete action space of four actions. In this variant, the agent’s time limit is reduced to 128 steps. Furthermore, MP can be made easier by including the path’s goal and origin in the agent’s observation. On the other hand, increasing the difficulty can be achieved by reducing the agent’s scale, decreasing its speed, or enlarging the path. In contrast to MM, episodes in MPGrid terminate earlier as the agent improves.
2.4 Endless Mystery Path
(a) Endless Path
(b) Agent Observation
Figure 5: In Endless Mystery Path, the path (a) is established by concatenating path segments, following a similar generation process as in MP. To avoid multiple neighboring path tiles, we add a transition tile (blue) between segments. Due to the constraints of representing an endless path within the observation space (b), we opt to visually fix the agent’s horizontal position. By rendering the path tiles behind the agent, the agent gains a sense of its horizontal motion, allowing it to perceive its progress along the path. Its horizontal position is adjustable, potentially obscuring more of the past path to increase the challenge.
Endless Mystery Path (EMP; Figure 5) introduces a never-ending path for the agent to traverse. Based on the original MP environment, we make several modifications to the path generation, the agent’s observation, and the terminal conditions.
The path generation in EMP (Figure 5(a)) involves concatenating path segments. Each segment is generated similar to MP, with the only difference being that the origin of a segment is based on the final tile of the previous segment. To ensure that path tiles do not have multiple neighbors, an additional tile is added between segments. The path is always generated from left to right. Consequently, we eliminate the agent’s ability to move left.
As the endless path cannot be fully captured in an $84\times 84$ RGB pixel observation, we visually fix the agent’s horizontal position. If nothing but the agent is shown in the observation, the agent lacks information about its horizontal motion. To address this, we render the path tiles behind the agent (Figure 5(a)), providing visual cues of the local horizontal position. For added difficulty, the visual horizontal position can be shifted further to the left, obscuring more information about the actual path.
Regarding terminal conditions, we aim to shorten episodes if the agent performs poorly. To achieve this, we implement three terminal conditions. Firstly, the agent is given a time budget of 20 steps to reach the next tile. Moving in the wrong direction or taking too long to progress within the budget leads to episode termination. The time budget is reset when the agent moves to the next tile or falls off and is relocated to the path’s origin. The second condition terminates the episode if the agent falls off before reaching its best progress. Lastly, the episode is terminated if the agent falls off at the same location for a second time. With these conditions in place, an episode can potentially last indefinitely if the agent consistently succeeds.
2.5 Searing Spotlights
[…]
(a) Annotated Ground Truth
(b) Agent Observation
(c) Visual Feedback
Figure 6: The environment’s annotated ground truth (a) reveals information about all relevant entities in Searing Spotlights. The top rows of pixels display the agent’s remaining health points, its last action, and indicate whether a positive reward was received during the last step. The last action is encoded using two chunks and three colors, while the last positive reward is represented by two colors. Yellow circles indicate coins, while the somewhat rounded gray shape represents the exit. When all coins are collected, the exit turns green. The floor of the environment is checkered blue and white. In the agent’s observation (b), the spotlights play a role in revealing or hiding other entities. As an additional form of visual feedback (c), the blue floor tiles turn red if a spotlight detects the presence of the agent.
Searing Spotlights (Figure 6) presents a pitch-black environment to the agent, where information is revealed only through roaming and threatening spotlights. The agent starts with a limited number of health points (e.g. five) and loses one health point per step if hit by a spotlight. The episode terminates if the agent runs out of health points or exceeds the time limit of 256 steps. To evade the imminent threats, the agent must strategically hide in the darkness. In order to escape from the encroaching spotlights, the agent needs to remember its past actions and at least one previous position to deduce its current location. This requires the agent to carefully manage its health point budget to ascertain its location, ideally with minimal exposure. To avoid monotonous behavior, two additional tasks are introduced in the environment, requiring the agent to actively traverse the environment. The first task asks the agent to collect a predetermined number of coins, while the second task involves reaching an exit to finish the episode once all the coins have been collected. Both the exit and the coins are randomly positioned during the environment’s reset phase. Collecting a coin rewards the agent with $+0.25$, while successfully utilizing the exit grants a reward of $+1$. By utilizing its memory effectively, the agent shall be able to deduce its current location and recall the positions of the coins and the exit.
A simplified grid version of Searing Spotlights is not available. As a default simplification, the episode starts with perfect information, and after a few steps, the global light is gradually dimmed until off. Just like in MP, more successful episodes terminate earlier.
2.6 Endless Searing Spotlights
Based on just a few modifications, we extrapolate the concept of Searing Spotlights to Endless Searing Spotlights (ESS). We eliminate the exit task and alter the coin collection task by spawning a new coin whenever the agent collects one. The newly spawned coin is always visible for six frames and the agent has a time budget of 160 steps to collect the new coin. Lastly, we raise the agent’s health points from five to ten.
3 Memory Baselines
In our experiments, we utilize the popular Deep Reinforcement Learning (DRL) algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). By default, PPO lacks a built-in mechanism to enable the ability of memory. To overcome this limitation, either recurrent neural networks (RNN) or self-attention mechanisms (i.e. transformer) are potential solutions. Prior works have extensively demonstrated the effectiveness of leveraging RNNs, such as Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Unit (GRU) (Cho et al., 2014), within the realm of DRL (Mnih et al., 2016; Espeholt et al., 2018; Huang et al., 2022). This also applies to previous studies employing transformers, like Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Hierarchical Chunk Attention Memory (HCAM) (Lampinen et al., 2021). Adopting these memory mechanisms is not trivially plug-and-play. They entail an increase in implementation complexity due to gathering more data, processing sequential data, and extending model interfaces. This may lead to a higher risk of errors and an undesired memory overhead in terms of hardware memory usage. To make these baselines more accessible to the community, we contribute two easy-to-follow baseline implementations of recurrent neural networks and Transformer-XL (TrXL) (Dai et al., 2019) in PPO.It is worth noting that there is a dearth of readily available and accessible transformer-based DRL implementations. Therefore, our realization of TrXL plus PPO is especially valuable in this context. In the following subsections, we provide an overview of the actor-critic model architecture that additionally leverages an auxiliary observation reconstruction loss. Finally, we will delve into the recurrent PPO and transformer-based PPO baselines.
3.1 General Model Architecture
Figure 7: Architecture of the actor-critic model featuring observation encoders and a memory encoder (green) that are shared among the policy, state-value function, and observation reconstruction heads.
Figure 7 presents a broad overview of the actor-critic model architecture. The encoding process begins with visual observations, and optionally, vector observations. To encode visual observations (i.e. pixel observations) of shape $84\times 84\times 3$, we employ the Atari CNN, a convolutional neural network that adheres to the standard topology originally developed for solving Atari games (Mnih et al., 2015). The vector observation is encoded by a multi-layer perceptron (MLP). The outputs of these encoders are concatenated and subsequently embedded into the input space of the memory encoder, which is based on either GRU or Transformer-XL. Once the data is propagated through the memory encoder, it is further processed by each individual head that may contain further hidden layers. The policy head (actor) samples multi-discrete actions as required by Memory Gym, while the value head (critic) approximates the state-value. Additionally, there is a head dedicated to reconstructing the original visual observation, leveraging the principles of self-supervised learning. Technically, the latent output of the memory encoder is fed to a transposed Atari CNN to achieve this reconstruction. The utility of such an auxiliary head is demonstrated by some of our results (Section 4) and prior works by Lampinen et al. (2021) and Hill et al. (2021). During development, we also tried to attach the reconstruction head directly to the encoder, but this led to catastrophic results.
3.2 Loss Function
Our baselines utilize PPO’s clipped surrogate objective (Schulman et al., 2017). Due to the usage of the memory encoder, the to be selected action $a_{t}$ of the policy $\pi_{\theta}$ depends on the current observation $o_{t}$ and the memory encoder’s output that we refer to as hidden state $h_{t}$. $\hat{A}_{t}$ denotes advantage estimates based on generalized advantage estimation (GAE) (Schulman et al., 2016), $\theta$ the parameters of a neural net and $\epsilon$ the clip range.
$L^{C}_{t}(\theta)=-\hat{\mathbb{E}}_{t}[\text{min}(q_{t}(\theta)\hat{A}_{t},\text{clip}(q_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t})]$ (5)
with $\textnormal{with ratio}~{}q_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|o_{t},h_{t})}{\pi_{\theta\text{old}}(a_{t}|o_{t},h_{t})}$
$t$ depicts the current time step. The value function is optimized using the squared-error loss $L_{t}^{V}(\theta)$. $\mathcal{H}[\pi_{\theta}](o_{t},h_{t})$ denotes an entropy bonus encouraging exploration (Schulman et al., 2017). The auxiliary observation reconstruction loss $L^{R}_{t}$ is based on the binary cross entropy. All three are weighted using the coefficients $c_{1}$, $c_{2}$, and $c_{3}$ and are added to $L_{t}^{C}$ to complete the loss:
$L^{CVHR}_{t}(\theta)=\hat{\mathbb{E}}_{t}[L^{C}_{t}(\theta)+c_{1}L^{V}_{t}(\theta)-c_{2}\mathcal{H}[\pi_{\theta}](o_{t},h_{t})+c_{3}L^{R}_{t}(o_{t},\hat{o}_{t})]$ (6)
$\textnormal{with}~{}L^{V}_{t}(\theta)=(V_{\theta}(o_{t},h_{t})-V_{t}^{target})^{2}$
$\textnormal{and}~{}L^{R}_{t}(o_{t},\hat{o}_{t})=-[\hat{o}_{t}\log(o_{t})+(1-\hat{o}_{t})\log(1-o_{t})]$
3.3 Recurrent Neural Network Baseline
(a) Inference
(b) Optimization
Figure 8: An overview of the distinct forward passes concerning inference and optimization when incorporating an RNN, such as GRU (stacked L times), as a memory encoder in a DRL model. During data sampling (8(a)), the agent interacts with its environment step-by-step, updating the prior hidden state based on encoded features of the current observation. In the optimization phase (8(b)), observations and formerly collected hidden states are divided into sequences. Each sequence is then processed by the RNN with its corresponding initial hidden state, followed by reshaping the resulting data back to its original batch dimensions.
In general, RNNs process sequential data iteratively by propagating information through recurrent connections. When utilizing these in DRL, the forward pass of the model differs between inference (sampling training data) and optimization as shown in Figure 8. During data sampling, the initial hidden state of the recurrent cell is initialized to zeros, and all output hidden states are stored for later optimization. During optimization, the sampled observations, actions, hidden states, and a loss mask are split into sequences based on a configurable sequence length. Sequences shorter than the designated length are zero-padded to optimize computational efficiency, while considering some memory overhead. Each sequence begins with its corresponding initial hidden state, enabling truncated back-propagation through time. Back-propagation through time, introduced by (Werbos, 1990), is an algorithm used to train recurrent neural networks by unfolding the network through time and applying the standard back-propagation algorithm (Rumelhart et al., 1986). Regarding the forward pass during optimization, the model is fed a batch of training data that is reshaped into sequences before being passed to the RNN. Once the data has propagated through the RNN, it is reshaped back to its original batch dimensions. Since only the RNN requires a sequence of data, it is more efficient to feed the non-recurrent components of the neural network the entire batch. Finally, during loss computation, the padded values are excluded using the loss mask, as defined below. $L^{mask}$ is the average over all losses not affected by the padding.
$L^{mask}(\theta)=\frac{\sum_{t}^{T}\left[mask_{t}\times L^{CVHR}_{t}(\theta)\right]}{\sum_{t}^{T}[mask_{t}]}$ (7)
$\textnormal{with}~{}mask_{t}=\begin{cases}0&\text{where padding is used}\\ 1&\text{where no padding is used}\\ \end{cases}$
Our final RNN baseline supports both GRU and LSTM architectures, along with truncated back-propagation through time. For our major experiments we chose GRU over LSTM for two main reasons. Firstly, GRU is computationally less expensive, as demonstrated in the work by Morad et al. (2023). Additionally, it is more GPU memory efficient, as it relies solely on its hidden state without the need for an additional cell state like how it is done in LSTM. Lastly, previous DRL studies showed that GRU slightly outperforms LSTM (Lambrechts et al., 2022; Morad et al., 2023). We also explored LSTM during hyperparameter tuning but observed inferior performance compared to GRU.
3.4 Transformer-XL Baseline
(a) Episodic Memory
(b) Transformer-XL Block
Figure 9: Figure 9(a) illustrates the episodic memory that stores past inputs to all Transformer-XL blocks for every step taken by the agent. Input sequences to the Transformer-XL blocks are retrieved using a sliding memory window. $T$ denotes the episode length. Figure 9(b) depicts the architecture of one Transformer-XL block that is stacked $B$ times. The dashed lines indicate that there is no flow of gradients into the episodic memory, the positional encoding, and the mask.
While RNNs process sequential data by propagating information through recurrent connections, transformers employ self-attention mechanisms to capture global dependencies in parallel across the entire sequence (Vaswani et al., 2017). Next, we detail our transformer baseline that draws conceptual inspiration from TrXL (Dai et al., 2019), GTrXL (Parisotto et al., 2020), and HCAM (Lampinen et al., 2021). The original transformer architecture proposed by Vaswani et al. (2017) is designed as a sequence-to-sequence model with an encoder-decoder structure. However, for the purpose of DRL, we adapt the architecture to a sequence-to-one model, focusing solely on the encoder as done in GTrXL and HCAM. This modification is driven by the underlying Markov Decision Process, which maps the environment’s current state to a single action rather than a sequence of actions.
To begin with, the input data can be conceptualized as a sequence of observations. In the context of visual observations, this is commonly referred to as frame stacking, where consecutive frames are combined to form the input sequence. However, instead of expensively stacking raw frames, we leverage the past outputs of the model’s embedding module (Figure 7), which serves as the input sequence to the first transformer block. The subsequent blocks make use of the past outputs from their preceding blocks. These sequences are stored in the agent’s episodic memory (Figure 9(a)) and are collected during inference based on the agent’s current timestep $t$. To accommodate computational limitations, we employ a sliding memory window approach, where the input sequence is derived from a fixed-length window. By leveraging past activations in the input sequence, the agent’s memory facilitates the ability to attend to events that extend beyond the boundaries of the fixed-length memory window, aligning with the segment-level recurrence mechanism of TrXL (Dai et al., 2019). Note that gradients are not back propagated through the episodic memory.
The architecture of our TrXL encoder block is depicted in Figure 9(b). At each timestep $t$, the block receives an input that is solely based on the current timestep. This input is added to the episodic memory and serves as the query $Q_{t}$ during the execution of multi-head attention (MHA) (Vaswani et al., 2017). The keys $K$ and values $V$ used in MHA are derived from the same data based on self-attention. Specifically, we retrieve $K$ and $V$ by slicing the episodic memory using the bounding indices $i=\max(0,t-\text{{window length}})$ and $t-1$. To ensure the positional information of $K$ and $V$ remains coherent, we add a positional encoding, following the approach from Vaswani et al. (2017). While learning a positional encoding can be effective, we found it to be less sample efficient. To restrict attention to timesteps up to the current timestep $t$, we employ a strictly lower triangular matrix as a mask in MHA. This mask is only necessary when $t$ is smaller than the length of the sliding memory window. The remaining configuration of the block adheres to the findings of Parisotto et al. (2020), which suggest improved performance with pre-layer normalization and the identity map reordering. We encountered vanishing gradient issues when applying layer normalization after MHA and the Position-Wise MLP, which is referred to as post-layer normalization. Although our implementation supports the gating mechanism of GTrXL, it resulted in either lower or equivalent performance while being computationally more expensive (Appendix C).
4 Experimental Analysis
To benchmark the effectiveness of our just introduced GRU and TrXL baselines, both powered by Proximal Policy Optimization (Schulman et al., 2017), we run empirical experiments on Memory Gym. Contrary to our previous study (Pleines et al., 2023), we exclude HELM (History comprEssion via Language Models) (Paischer et al., 2022a) and its successor, HELMv2 (Paischer et al., 2022b). The decision stems from HELM’s suboptimal wall-time efficiency, which is six times more expensive than our GRU agent, making hyperparameter optimization impractical for our purposes. As we are strongly limited by our compute budget of 25,000 GPU hours, we rather dedicated the resources towards the development of our endless environments and TrXL baseline. Section 5.2 discusses further reasons for not considering more baselines in our experiments. We also conduct preliminary tests using Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) on the finite environments without any hyperparameter tuning. These models usually underperformed when applied directly, suggesting that out-of-the-box applications might necessitate hyperparameter optimization.
As part of our evaluation protocol, the following sample efficiency curves depict the interquartile mean (IQM) along with a stratified bootstrapped confidence interval of 0.95 (shaded area on the plots), following the guidelines of Agarwal et al. (2021). Each training run is repeated 5 times. To evaluate the generalization ability of our baselines, we extensively test them on 50 novel environment seeds that were excluded from the training process. This evaluation is repeated 3 times for each data point, resulting in a comprehensive assessment based on a total of 750 episodes for every data point presented.
The next subsection presents a two-fold examination of the finite environments. Firstly, we reaffirm that memory is imperative in simpler environment instances, a conclusion previously drawn in our earlier study (Pleines et al., 2023). Enhanced hyperparameters allow us to report improved results. Secondly, we compare the performance of GRU and TrXL on the original finite tasks. TrXL exhibits stronger policies in Mortar Mayhem than GRU, while it is more sample efficient in Mystery Path. Searing Spotlights is an exception to this where TrXL is less sample efficient. Contradicting our initial expectations and the results on the finite environments, we unveil surprising findings that highlight GRU’s superiority over TrXL in all endless environments. For a comprehensive overview of the hyperparameters employed in our experiments, see Appendix B. Results for GTrXL and LSTM on the finite environments can be found in Appendix C. Additionally, a selection of wall-time efficiency statistics is available in Appendix D.
4.1 Finite Environments: Transformer-XL is often more efficient than GRU
(a) Mortar Mayhem Act Grid
(b) Mortar Mayhem Grid
(c) Mortar Mayhem
(d) Mystery Path Grid (Cues On)
(e) Mystery Path Grid
(f) Mystery Path
(g) Searing Spotlights
Figure 10: Performance comparison across instances of Memory Gym’s finite environments. Mortar Mayhem Act Grid (a) provides commands as a fully observable vector, bypassing the Clue Task. Both (a) and Mortar Mayhem Grid (b) utilize grid-like locomotion. (c) represents the default Mortar Mayhem task. Mystery Path Grid operates on grid-like locomotion, with (a) rendering the the goal and origin to the agent’s observation. (b) and (c) hide these, while (c) is the default Mystery Path task. Searing Spotlights (g) is not varied, while all conducted runs leverage observation reconstruction (Obs. Rec.) in this environment.
Before assessing GRU and TrXL on the default challenges of the finite environments, the first set of experiments focuses on confirming that the agents trained on Memory Gym indeed require memory. To accomplish this, we train several minor baselines alongside GRU and TrXL on simplified versions of the environments. These baselines include PPO without memory, PPO with frame stacking (4 and 16 grayscale frames), and PPO with relative positional encoding (Vaswani et al., 2017) provided as vector observation.
Results on Mortar Mayhem (Figures 1010(a)-10(c)) show the number of commands that were properly executed during one episode. As mentioned in section 2.1, the agent is not required to memorize the command sequence in Mortar Mayhem Act Grid as it is included in the agent’s vector observation. Under these simplified conditions, PPO without memory is ineffective, while the frame stacking agents achieve nearly either 5 or 8 commands with a slight upward trend. Agents using GRU or TrXL are the most effective, completing all 10 commands, while those with relative positional encoding follow closely, completing 9.5 commands. The positional encoding baseline adds the episode’s current time step to the agent’s observation allowing for temporal awareness within the sequential execution of commands. Once the Clue Task is present (Figure 10(b)), only GRU and TrXL prove to be effective and therefore we conclude that Mortar Mayhem demands memory.
We obtain a similar impression when training on Mystery Path Grid where the origin and goal are not perceivable by the agent (Figure 10(e)). In this case, TrXL outperforms GRU in terms of sample efficiency, while both methods are effective. If the origin and the goal are visible, a horizon of 16 frames is sufficient to train an effective policy as shown by the frame stacking agent (Figure 10(d)). Stacking 4 frames or leveraging positional encoding leads to an IQM success rate of 40% with a slight trend upward. The memory-less agent’s success rate goes up to about 11%.
No success is proven by the minor baselines in Searing Spotlights as seen in Figure 10(g). Only in Searing Spotlights, the minor baselines also utilize observation reconstruction. Utilizing the observation reconstruction loss significantly improves the sample efficiency of GRU and TrXL. The reason for this can be found in rare events: the agent, the coin, and the exit are seldomly visible. Hence, the visual encoder takes advantage from the auxiliary learning signal provided by the observation reconstruction. However, both GRU variants, with and without observation reconstruction, exhibit notably higher sample efficiency compared to their respective TrXL counterparts.
When facing the default task of Mortar Mayhem, a different results becomes apparent (Figure 10(c)). TrXL with observation reconstruction solves all 10 commands after training for about 100 million steps, while GRU converges to only about 6 commands. GRU’s performance drops to about 5 commands if the observation reconstruction loss is used. If TrXL trains without this auxiliary loss, its performance drops to about 2 successful commands.
In the case of Mystery Path (Figure 10(f)), we opted not to train agents using observation reconstruction, as the always visible agent is the only entity that requires visual encoding. In this scenario, TrXL achieves the desired success rate in approximately 100 million steps, whereas GRU requires twice as many steps to achieve the same level of performance.
In summary, TrXL proves to be more effective than GRU in Mortar Mayhem and exhibits higher sample efficiency in the case of Mystery Path. However, it is worth noting that GRU, both with and without observation reconstruction, outperforms its respective TrXL counterparts in Searing Spotlights.
4.2 Endless Environments: GRU is more effective than Transformer-XL
(a) Endless Mortar Mayhem
(b) Endless Mystery Path
(c) Endless Searing Spotlights
Figure 11: GRU’s consistent superiority over TrXL is revealed in Memory Gym’s endless environments. The less opaque line marked with crosses depicts the interquartile mean (IQM) episode length that refers to the right Y-Axis.
Across all three endless environments, the recurrent agent (GRU) consistently outperforms the transformer-based agent (TrXL) by a significant margin. Notably, the most substantial performance gap between GRU and TrXL emerges in the results of Endless Mortar Mayhem (Figure 11(a)). Within this context, GRU attains an impressive IQM of 120 executed commands using roughly 300 million steps, whereas TrXL only attains an IQM of 18 commands. Furthermore, the incorporation of observation reconstruction exacerbates this performance difference, leading TrXL to a further deterioration, resulting in an IQM of 10. Furthermore, GRU exhibits consistently longer IQM episode lengths, with TrXL achieving an IQM of 288 steps and GRU reaching a peak IQM episode length of 3462 steps.
The outcomes depicted in Figure 11(b) illustrate the results obtained from the Endless Mystery Path environment. In this case, GRU remains more effective than TrXL, although the gap between their performances is narrower compared to EMM. GRU achieves an IQM return of 2.65, with TrXL achieving 2.02. Although neither memory approach converges within the allocated training budget of 819 million steps, GRU displays a more pronounced upward trend in performance. In terms of episode length, neither agent achieves lengths as seen in EMM. GRU attains an IQM episode length of 600 steps in EMP, while TrXL reaches up to 400 steps. This divergence can be attributed to the episode’s dynamics, where agents tend to fall off after significant progress, necessitating more steps to regain lost ground and accumulate further rewards.
In Endless Searing Spotlights (Figure 11(c)), GRU is about 3.3 times more effective than TrXL. Both agents utilize the observation reconstruction loss, but TrXL converges to an IQM of 6 collected coins after approximately 200 million steps, while GRU achieves an IQM of about 20 coins, continuing to show an upward trend in its performance. The IQM episode length for TrXL is 200 steps, while GRU achieves a longer IQM episode length of 600 steps.
4.3 Investigating Transformer-XL’s Surprisingly Low Effectiveness
Figure 12: Experiments on Endless Mortar Mayhem with varied TrXL configurations. The learning rate schedule (LR) is adjusted to decay from $42.75e-$ 4 to $41.0e-$ 4 over 160 million steps, compared to the previous $51.0e-$ 5. We also incorporate observation reconstruction (Obs. Rec.) and observe the most significant improvement upon introducing a ground truth estimation head (GT) as sanity check. Another test augments TrXL’s query with relative positional encoding (QPos). The final, but inferior, experiment combines the optimized learning rate, augmented query, and ground truth estimation.
The unexpectedly low performance of TrXL prompts us to several hypothesis that are explored within the next subsections. Endless Mortar Mayhem poses the largest performance gap and therefore we utilize this environment for further investigations.
4.3.1 Inadequate Networks Capacity
Our TrXL baseline comprises 2.8 million trainable parameters, while GRU consists of 4.05 million, prompting the question of whether TrXL’s model architecture lacks the necessary capacity. To investigate this, we conducted experiments varying the number of blocks (2, 3, 4), the embedding dimension size (256, 384, 512), and the memory window length (256, 384, 512). None of these adjustments led to closing the performance gap to GRU.
Scaling up TrXL also increases its demand for GPU memory. When scaling up multiple architecture details, our implementation and hyperparameters may cause the training to exceed the available 40GB GPU memory of an NVIDIA A100. Workarounds include reducing the batch size or transferring training data between the GPU and CPU, but these options significantly worsen wall-time efficiency.
Furthermore, another indication of untapped capacity lies in the amplification of the learning signal, which we elaborate on in the following section.
4.3.2 Weak Learning Signal
Figure 12 depicts several additional experiments aimed at amplifying the learning signal. We made naive a adjustment to the learning rate schedule, utilized observation reconstruction, and introduced a ground truth estimation head. This ground truth estimation head is added to the actor-critic model architecture and predicts the target position to move to. Labels are provided by the environment, and the estimates are optimized using the mean-squared error loss. While ground truth information is typically unavailable and considered inappropriate for this benchmark, it serves as a useful sanity check. It helps determine if an additional learning signal is advantageous in this context. When training incorporates ground truth estimation, there is a noticeable improvement in the agent’s policy. Previously, the IQM of completed commands stood at 18; with ground truth estimation, it reaches 50. This outcome implies that the model’s capacity might not be the primary constraint. However, using the ground truth estimation head can lead to instability in training, with the agent’s performance fluctuating between 20 and 50 completed commands after about 300 million steps.
During the initial 160 million training steps, the learning rate linearly decays from $2.75$e-$4$ to $1.0$e-$5$. By setting the final decayed learning rate to $1.0$e-$4$, the IQM performance hits 24 commands. Lower performances are observed when adding the observation reconstruction head to either just TrXL or TrXL with the optimized learning rate schedule.
In summary, augmenting the learning signal — either by adjusting the learning rate or by incorporating ground truth estimation — leads to more performant policies.
4.3.3 Lack of Temporal Information in the Initial Query
The next hypothesis pertains to the initial query of the first TrXL block. This query is constructed solely based on the features extracted from the observation encoders at the current time step, encompassing only minimal temporal information. In contrast, the query of the second block draws from the aggregated outcome of the memory window, thus capturing more substantial temporal information. However, we believe that the initial query could be further enriched with additional information. In Figure 12, one experiment augments the query with relative positional encoding, providing direct access to the current time step. Despite this enhancement, the agent’s performance only moderately improves, reaching an IQM of 19 completed commands. We also explored embedding the original query with the previous output of the last TrXL block. However, this simple approach did not yield positive results.
All of the aforementioned measures by far do not reach the level of the GRU agent. Even if combined, a low performance of 33 executed commands is achieved, which is also relatively unstable.
4.3.4 Harmful Off-Policy Data in the Episodic Memory
At last it can be discussed whether the combination of PPO and TrXL’s episodic memory pose a threat. PPO is an on-policy algorithm that consumes the training data for a few epochs. Once the second training epoch on the same batch commences, the batch is already considered off-policy. However, PPO’s loss function can mitigate this issue and hence allows for several epochs. But this mitigation is likely limited and we wonder whether this issue is enlarged by the episodic memory. In our current training setup, an agent collects 512 steps of data. If an episode exceeds 512 steps and is still running, the earlier steps from that episode become outdated over several epochs. In future work, it needs to be investigated if a stale episodic memory, which is based on off-policy data, hinders training progress. Frequently refreshing the episodic memory is expected to be expensive in terms of wall-time.
5 Related Work
We begin by describing different approaches aimed at improving agents’ capabilities through memory mechanisms, and highlight the challenge of accessing related baselines. Subsequently, we present an overview of existing memory benchmarks, comparing them conceptually to Memory Gym.
5.1 Memory-based Approaches
The work of Mnih et al. (2015) on mastering Atari games (Bellemare et al., 2013) using Deep Q-Networks (DQN) marked a significant milestone in the advancement of deep reinforcement learning algorithms. To capture temporal information, such as velocities and accelerations, a frame stacking approach was employed. More advanced memory-based approaches were introduced, which we will briefly detail subsequently.
Deep Recurrent Q-Networks (DRQN) (Hausknecht and Stone, 2015) combine DQN with an LSTM. The replay buffer stores sequences, while initializing the recurrent hidden state to zero during training. The effectiveness of this approach is demonstrated on Atari games that involve frame flickering.
Recurrent Replay Distributed DQN (R2D2) (Kapturowski et al., 2019) extends DRQN with a distributed prioritized experience replay buffer, storing overlapping sequences of fixed length. A burn-in period is utilized to address staleness in recurrent hidden states by propagating a portion of the sequence for a more effective starting state. R2D2’s performance is evaluated on Atari games (Bellemare et al., 2013) and DMLab-30 (Beattie et al., 2016).
The Memory Recall Agent (MRA) (Fortunato et al., 2019), built on IMPALA (Espeholt et al., 2018), employs a two-fold memory mechanism: a working memory based on an LSTM and a slot-based episodic memory containing summaries of past experiences. By utilizing a query-key approach, information is read from and written to the episodic memory. MRA is evaluated on the Memory Task Suite contributed by the same publication.
Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) adds the concept of Transformer-XL (TrXL) to enable an agent, trained with V-MPO (Song et al., 2020), to leverage memory. By incorporating an identity map reordering and a gating mechanism inspired by GRU, TrXL is further improved. GTrXL’s performance is demonstrated on DMLab-30 (Beattie et al., 2016), Numpad, and Memory Maze.
The Hierarchical Chunk Attention Mechanism (HCAM) (Lampinen et al., 2021) builds upon TrXL by partitioning the agent’s memory into chunks and selecting the most relevant chunks based on their summaries to attend to their stored information in detail. HCAM is trained using IMPALA (Espeholt et al., 2018) and evaluated on multiple environments including Dancing the Ballet, Object Permanence, Rapid Word Learning, Visual Match, Paired Associative Inference, and One-Shot StreetLearn Navigation.
History comprEssion via Language Models (HELM) (Paischer et al., 2022a) utilizes pre-trained language models to compress a history of observations into a single context vector (memory). It employs a frozen TrXL encoder, pre-trained on the WikiText 103 dataset (Merity et al., 2017). HELM is evaluated on the Procgen (Cobbe et al., 2020) and Minigrid (Chevalier-Boisvert et al., 2018) environments.
5.2 Lack of Accessible Memory Baselines
The accessibility of the approaches to memory described above poses a common issue. While HELM is fully open source and successfully integrated into our training framework, the implementations of DRQN, R2D2, MRA, and GTrXL have not been made publicly available by their original authors. HCAM is partially open source regarding its model architecture. The lack of open source implementations makes reproducing their results more challenging, as implementation details play a crucial role. This issue is particularly pronounced for GTrXL, as to the best of our knowledge, no one has yet been able to reproduce its underlying DRL algorithm, V-MPO.
Within the community, RLlib (Liang et al., 2018), DI-engine (DI-engine Contributors, 2021), and Brain Agent (Lee et al., 2022) advertise to feature GTrXL. However, these frameworks are large and complex, supporting multiple DRL algorithms, model architectures, and environments, which hampers verification, debugging, gaining insights, and making modifications. We tested RLlib with Memory Gym, but obtained negative returns even though the environments do not signal negative rewards. DI-Engine’s GTrXL implementation is based on R2D2, and our experiments revealed poor sample throughput by a magnitude of 10 compared to our GRU baseline. When debugging Brain Agent’s implementation, we noticed an unexpected discrepancy in the query length between data sampling and optimization. While we do not claim that these frameworks are dysfunctional, we present our limited experience in exploring them. We believe that our contributed open source baselines are easy to follow, which facilitates reproducibility and further research in the field.
5.3 Related Reinforcement Learning Memory Benchmarks
Previous studies have explored the use of memory-based agents in a variety of environments. Some of them are originally fully observable but are turned into partially observable Markov Decision Processes (POMDP) by adding noise or masking out information from the agent’s observation space. For instance, this was done for the Arcade Learning Environment (Bellemare et al., 2013) by using flickering frames (Hausknecht and Stone, 2015) and common control tasks by removing the velocity from the agent’s observation (Heess et al., 2015; Meng et al., 2021; Shang et al., 2021).
Control tasks also touch on the context of Meta-Reinforcement Learning, where memory mechanisms are prominent (Wang et al., 2021; Melo, 2022; Ni et al., 2022). The same applies to Multi-Agent Reinforcement Learning (Berner et al., 2019; Baker et al., 2020; Vinyals et al., 2019). As we solely focus on benchmarking the agent’s memory and its ability to generalize, we do not compare Memory Gym to environments of more complex contexts such as DM Alchemy (Wang et al., 2021), Crafter (Hafner, 2021), or Obstacle Tower (Juliani et al., 2019). These might need additional components to the agent’s architecture and its training paradigm, for instance to approach a notable exploration challenge.
When skimming the approaches described in section 5.1, it appears that there is a lack of consensus regarding the choice of environments for evaluation as they assess their approaches on different benchmarks, thereby hindering comparisons between them. Reasons for this observation are either missing accessibility, slow simulation speeds, or weak memory challenges (Pleines et al., 2023). Subsequently we give a coarse overview of recently used benchmarks chronologically.
Deepmind Lab 30 (Beattie et al., 2016) presents a collection of 30 procedurally generated first-person 3D environments. Starting from a general motor-control navigation task and visual observations, each environment poses a different goal such as collecting fruit, playing laser tag or traversing a maze. Parisotto et al. (2020) categorize these environments into memory and reactive tasks.
Minigrid (Chevalier-Boisvert et al., 2018) includes a 2D grid memory task inspired by the T-Maze (Wierstra et al., 2007). In this environment, the agent is required to memorize an initial goal cue, traverse a long alley, and correctly choose the exit once the end is reached.
Miniworld (Chevalier-Boisvert, 2018) encompasses a set of first-person 3D environments based on the tasks introduced by Minigrid.
VizDoom (Wydmuch et al., 2018) features first-person shooter 3D environments that revolve around motor-control navigation tasks, with the added challenge of computer-controlled enemies that may confront the agent.
The Memory Task Suite (Fortunato et al., 2019) includes diverse memory-based environments across four categories: PsychLab (Leibo et al., 2018) for image-based tasks (e.g. detect change), Spot the Difference for cue memorization, goal navigation tasks inspired by the Morris water maze (D’Hooge and De Deyn, 2001), and transitive object ordering tasks.
Procgen (Cobbe et al., 2020) consists of 16 2D environments that offer procedurally generated and fully observable levels. These environments encompass a combination of puzzle-solving, platforming, and action-based games. 6 of these environments have been modified to become partially observable by reducing and centering the agent’s visual observation.
Numpad (Parisotto et al., 2020) requires the agent to accurately press a series of keys in a specific sequence of fixed length. The sequence itself is not provided to the agent, resulting in the agent employing a trial-and-error approach while memorizing the underlying sequence.
Memory Maze (Parisotto et al., 2020) shares similarities with the Morris Water maze (D’Hooge and De Deyn, 2001), as the agent must locate an apple, then undergoes random repositioning, requiring it to find the apple’s location again and again.
Dancing the Ballet (Ballet) (Lampinen et al., 2021) presents a sequence of up to 8 dances visually, with the agent remaining stationary. After all the dances are displayed, the agent is required to identify a specific one to successfully complete the task.
PopGym (Morad et al., 2023) has been published at ICLR 2023 like our original work contributing Memory Gym (Pleines et al., 2023). It is a benchmark consisting of 15 environments that are tagged as diagnostic, control, noisy, game, or navigation. These environments are designed to quickly converge by providing vector observations instead of visual ones.
5.4 Comparing Memory Gym to Related Benchmarks
Memory Gym’s environments offer a unique feature that sets them apart from related environments: an endless design tailored to rigorously test memory capabilities. It is crucial to differentiate between “endless” in this context and “open-ended” environments, such as Minecraft (Fan et al., 2022). While open-ended environments aim for agents to autonomously acquire a broad range of diverse and general skills without human oversight, Memory Gym’s ”endless” design remains anchored to a specific task structure.
The finite environments Mortar Mayhem and Mystery Path exhibit both similarities and unique characteristics compared to related environments. Several environments, such as Ballet, PopGym’s Autoencode, Spot the Difference, and Minigrid Memory, require the agent to initially memorize at least one cue to solve the task. While Ballet, Spot the Difference, and Minigrid Memory involve mapping clues to a single solution, Mortar Mayhem and PopGym’s Autoencode challenge the agent to solve a sequential puzzle. In Mortar Mayhem, delays between showing cues and executing commands can be dynamically sampled, similar to the concept in Ballet. Mystery Path is most similar to the navigation concepts of the environments that are inspired by the Morris water maze. Once fallen off, the agent must return to the point before where it lastly fell off and then has to take different actions to make progress on the hidden path. Concerning Searing Spotlights we were unable to identify an already existing environment that shares conceptual details.
6 Conclusion
In this study, we advanced Memory Gym’s environments to novel endless tasks, tailored to benchmark memory-based Deep Reinforcement Learning algorithms. These tasks feature a mounting challenge similar to the car game “I packed my bag”. As the agent’s policy refines, the task dynamically expands, serving as an automatic curriculum. This innovative framework enables a thorough evaluation of memory-based agents, emphasizing their overall effectiveness beyond mere interaction efficiency with the environment.
To foster a more inclusive research landscape, we contributed an open-source PPO baseline powered by Transformer-XL (TrXL). This baseline employs an attention mechanism applied to an episodic memory with a sliding window. When benchmarked against Memory Gym, the TrXL baseline competes head-to-head with another prominent baseline rooted in Gated Recurrent Unit (GRU). Notably, in finite environments, TrXL demonstrates superior effectiveness in Mortar Mayhem and enhanced sample efficiency in Mystery Path. However, in the Searing Spotlights environment, GRU emerges as the more efficient architecture.
Perhaps our most unexpected revelation is the comeback of GRU in the endless environments. GRU surpassed TrXL by large margins, while also being computationally more efficient. Further probing into the Endless Mortar Mayhem environment revealed only marginal improvements for TrXL upon enriching the learning signal.
This prompts future investigations into potential reasons behind TrXL’s limitations, such as the possibility of episodic memory staleness or an initial query lacking temporal awareness. As we move forward, it will be compelling to discern the performance thresholds of other memory mechanisms and DRL algorithms. In addition to examining more attention-based and recurrent architectures, it may be worth exploring structured state space sequence models (Gu et al., 2022), given their recent debut in DRL (Lu et al., 2023).
Yet, the research community still grapples with the absence of a standardized benchmark. While we do not consider Memory Gym to satisfy this role, we believe it to be an invaluable complement to benchmarking memory-based agents. Addressing this issue necessitates a thorough deliberation on the essential criteria that a complete memory benchmark should encompass. For instance, should an ideal memory benchmark incorporate environments with varied modalities in their observation and action spaces?
Acknowledgements
Our sincere appreciation goes to Vincent-Pierre Berges for his enlightening discussions that greatly enriched our work. We also extend our gratitude to Andrew Lampinen for his invaluable insights on utilizing Transformer-XL as episodic memory. Special thanks are due to Günter Rudolph for his unwavering support. This project would not have been possible without the generous computing time provided by the Paderborn Center for Parallel Computing (PC2) and the support from the Linux-HPC-Cluster (LiDO3) at TU Dortmund. We are deeply thankful for all of their contributions.
Appendix A Environment Parameters
Table 1: Default reset parameters of the finite environments. Parameters marked with an asterisk (*) indicate uniform sampling. Values enclosed in square brackets represent discrete choices, while values in parentheses denote a range from which sampling is performed. For Mortar Mayhem Grid and Mystery Path Grid, the parameters remain the same as its parent, with only the modified ones presented.
Mortar Mayhem Searing Spotlights
Parameter Default Parameter Default
Agent Scale 0.25 Max Episode Length 256
Agent Speed 3 Agent Scale 0.25
Arena Size 5 Agent Speed 3
No. Available Commands 9 Agent Always Visible False
No. Commands* [10] Agent Health 5
Command Show Duration* [3] Sample Agent Position True
Command Show Delay* [1] Use Exit True
Execution Duration* [6] Exit Scale 0.5
Execution Delay* [18] Exit Visible False
Show Visual Feedback True Number of Coins* [1]
Reward Command Failure 0 Coin Scale 0.375
Reward Command Success 0.1 Coin Always Visible False
Reward Episode Success 0 No. Initial Spotlight Spawns 4
Mortar Mayhem Grid No. Spotlight Spawns 30
Parameter Default Spotlight Spawn Interval 30
No. Available Commands 5 Spotlight Spawn Decay 0.95
Execution Duration* [2] Spotlight Spawn Threshold 10
Execution Delay* [6] Spotlight Radius* (7.5-13.75)
Mystery Path Spotlight Speed* (0.0025-0.0075)
Parameter Default Spotlight Damage 1
Max Episode Length 512 Light Dim Off Duration 6
Agent Scale 0.25 Light Threshold 255
Agent Speed 3 Show Visual Feedback True
Cardinal Origin Choice* [0, 1, 2, 3] Show Last Action True
Show Origin False Show Last Positive Reward True
Show Goal False Render Background Black False
Show Visual Feedback True Hide Checkered Background False
Reward Goal 1 Reward Inside Spotlight 0
Reward Fall Off 0 Reward Outside Spotlights 0
Reward Path Progress 0.1 Reward Death 0
Reward Step 0 Reward Exit 1
Mystery Path Grid Reward Coin 0.25
Parameter Default Reward Max Steps 0
Max Episode Length 128
Reward Path Progress 0
Table 2: Default reset parameters of the endless environments. Parameters marked with an asterisk (*) indicate uniform sampling. Values enclosed in square brackets represent discrete choices, while values in parentheses denote a range from which sampling is performed. If the max episode length is equal or less than zero, episodes will not terminate because of reaching the max episode length.
Endless Mortar Mayhem Endless Searing Spotlights
Parameter Default Parameter Default
Max Episode Length -1 Max Episode Length -1
Agent Scale 0.25 Agent Scale 0.25
Agent Speed 3 Agent Speed 3
No. Available Commands 9 Agent Always Visible False
Command Show Duration* [3] Agent Health 10
Command Show Delay* [1] Sample Agent Position True
Execution Duration* [6] Coin Scale 0.375
Execution Delay* [18] Coin Show Duration 6
Show Visual Feedback True Coin Always Visible False
Reward Command Failure 0 Steps per Coin 160
Reward Command Success 0.1 No. Initial Spotlight Spawns 3
Endless Mystery Path Spotlight Spawn Interval 50
Parameter Default Spotlight Radius* (7.5-13.75)
Max Episode Length -1 Spotlight Speed* (0.0025-0.0075)
Agent Scale 0.25 Spotlight Damage 1
Agent Speed 3 Light Dim Off Duration 6
Show Origin False Light Threshold 255
Show Past Path True Show Visual Feedback True
Show Background False Render Background Black False
Show Stamina False Hide Checkered Background False
Show Visual Feedback True Show Last Action True
Camera Offset Scale 5 Show Last Positive Reward True
Stamina Level 20 Reward Inside Spotlight 0
Reward Fall Off 0 Reward Outside Spotlights 0
Reward Path Progress 0.1 Reward Death 0
Reward Step 0 Reward Coin 0.25
Appendix B Hyperparameters
Table 3: Hyperparameters and architectural details used in our final experiments. The “Final” column denotes the selected hyperparameter values, while the “Search Space” column represents additional discrete choices explored during the tuning process. The last column details the old hyperparameters of our previous study (Pleines et al., 2023). For this column alone, empty cells indicate that the values align with those in the “Final” column. T-Fixup is a transformer weight initialization approach by Huang et al. (2020).
Hyperparameter Final Search Space Old
Training Seeds 100000
Number of Workers 32
Worker Steps 512
Batch Size 16384
Disocunt Factor Gamma 0.995 0.99
GAE Lamda 0.95
Optimizer AdamW
Epochs 3 2 4
Number of Mini Batches 8 4
Advantage Normalization No Batch Mini Batch Mini Batch
Clip Range Epsilon 0.1 0.2 0.3 0.2
Value Loss Coefficient 0.5 0.25 0.25
Initial Learning Rate 2.75e-4 2.0e-4 3.0e-4 3.5e-4 3.0e-4
Final Learning Rate 1.0e-5 1.0e-4
Initial Entropy Coef. 1.0e-4 1.0e-3 1.0e-2
Final Entropy Coef. 1.0e-6 1.0e-5
Reconstruction Loss Coef. 0.1 0.5 1.0 n/a
Maximum Gradient Norm 0.25 0.35 0.5 1.0 0.5
Recurrent Neural Network
Number of Recurrent Layers 1 2
Embedding Layer Yes No
Layer Type GRU LSTM
Residual False True
Sequence Length 512 or max
Hidden State Size 512 256 384
Transformer-XL
Number of TrXL Blocks 3 2 4
Block Dimension 384 256 512
Block Weight Initialization Xavier Orthogonal Kaiming T-Fixup
Positional Encoding Relative None Learned
Number of Attention Heads 4 8
Memory Window Length 256 or max
Table 3 presents the hyperparameters utilized in our final experiments unless otherwise specified. The learning rate and entropy coefficient undergo linear decay from their initial to final values. This decay occurs exclusively during the first 10,000 PPO updates (equivalent to 163,840,000 steps). As training progresses beyond this point, the final hyperparameter serves as a lower threshold. Regarding the sequence length and memory window length, they are determined based on the longest possible episode in the finite environments. In the case of endless environments, the so far best sequence length is fixed at 512 for GRU and 256 for TrXL.
Our hyperparameter search was conducted using a customized implementation built upon the optuna tuning framework (Akiba et al., 2019). The objective was to discover a set of hyperparameters that demonstrate strong performance across both our GRU and TrXL baselines and Memory Gym’s environments. Considering limited resources and the computational expense of tuning, we restricted the search space to a limited number of options. The majority of tuning experiments focused on the finite environments. Therefore, we do not claim to provide optimal hyperparameters.
For each environment, we established individual tuning experiments corresponding to the choices presented in Table 3. Each experiment, or trial, involved a baseline set of hyperparameters that was modified by selecting a single parameter choice from the entire search space. Notably, we did not train permutations of all available choices at this stage. To optimize resource utilization, we implemented optuna’s percentile pruner, which retained the top 25 percentile of trials. Trials that performed below this percentile were pruned, but only after surpassing 2,000 PPO updates as a warm-up phase.
After completing the single experiments, we expanded the hyperparameter search by allowing for permutations sampled using optuna’s Tree-structured Parzen Estimator. It is important to note that individual trials were not repeated, but since tuning was conducted across multiple environments, one could consider them as repetitions in a broader sense.
The batch size is determined by multiplying the number of environment workers by the number of worker steps performed by each worker to gather training data. To efficiently utilize the resources of a system with 32 CPU cores and 40GB VRAM on an nVidia A100 GPU, we chose a batch size of 16,384 samples. However, when scaling up the model, it is possible to exceed the available GPU memory. In such cases, there are two options: reducing the batch size or outsourcing the data to CPU memory. Both of these workarounds come with increased expenses in terms of wall-time.
Appendix C GTrXL and LSTM Results on the Finite Environments
(a) Mortar Mayhem
(b) Mystery Path
(c) Searing Spotlights
Figure 13: GTrXL and LSTM results on the finite environments.
Figure 13 presents extended results for the finite environments. The performance metrics for TrXL and GRU agents remain consistent with those depicted in Figure 10. However, this figure displays twice as many data points on the x-axis. We include results from agents utilizing Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and the Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997). These models are benchmarked without any hyperparameter tuning. For GTrXL, biases of both 0 and 2 are tested, based on the suggestion by Parisotto et al. (2020) that a larger bias might enhance learning. All additional agents in the Searing Spotlights environment employ the observation reconstruction loss (Obs. Rec.), while it is omitted in the Mystery Path environment.
For Mortar Mayhem (as seen in Figure 13(a)), the LSTM agent settles at 3 commands, whereas the GTrXL agents using observation reconstruction both achieve an IQM of 2 commands. In the Mystery Path results (Figure 13(b)), the LSTM agent achieves full success, albeit requiring more samples compared to the GRU and TrXL agents. The GTrXL agent with a bias of 0 reaches a 94% success rate, while the one with a bias of 2 achieves only 81%, challenging the earlier assertion by Parisotto et al. (2020). The confidence interval for these results indicates potential training instability. In the Searing Spotlights environment, the LSTM agent with observation reconstruction is slightly less efficient than its GRU counterpart. However, GTrXL agents, though successful, lag behind TrXL with observation reconstruction.
It is important to note that we avoid making conclusive statements based on these results. This caution stems from the fact that the LSTM and GTrXL baselines are not tuned, and the GTrXL implementation awaits further verification.
Appendix D Wall-Time Efficiency
Table 4: Wall-time efficiency metrics for various agents trained on Endless Mortar Mayhem. Agents are arranged from left to right based on increasing mean training duration, measured in hours. Other values are represented in seconds and comprise the time needed for a single update (i.e. one complete iteration of running PPO). The second row indicates if the neural network was compiled before training. Note: TrXL* metrics use 4k initial training data points, while other runs utilize 150k data points covering the full training process.
GRU TrXL
TrXL +
QPos
TrXL +
Obs. Rec.
TrXL +
LR + QPos + GT
TrXL*
Compiled No Yes Yes Yes Yes No
Mean 4.11 5.53 5.6 5.76 5.86 7.97
Std 0.35 1.12 1.31 1.1 1.68 0.25
Min 3.6 5.2 5 5.4 5.3 7.6
Median 4.1 5.4 5.5 5.7 5.8 7.9
IQM 4.03 5.48 5.53 5.72 5.82 7.93
Mean Duration 57.08 76.81 77.78 80 81.39 110.69
Reliably reporting the wall-time of all conducted experiments is not feasible due to varying circumstances as varying hardware or different loads of the file system. Therefore, we only provide a selection of measurements on Endless Mortar Mayhem and Mystery Path.
Table 4 presents wall-time efficiency metrics for various agents trained on Endless Mortar Mayhem, utilizing the noctua2 high-performance cluster [^2] https://pc2.uni-paderborn.de/de/hpc-services/available-systems/noctua2. We leverage 32 CPU cores (AMD Milan 7763), an NVIDIA A100 GPU, and 100GB RAM for one training run. With PyTorch version 2, models can be lazily compiled at training onset, reducing TrXL’s training time by a significant 30%. Previously, plain TrXL training averaged 110 hours, but this has been reduced to 76 hours. Incorporating observation reconstruction adds roughly 3 hours, and with ground truth estimation, it extends to 81 hours. The GRU agents are the most wall-time efficient, completing in approximately 57 hours. Thus, recurrence emerges as the most effective and efficient architecture in both sample and wall-time metrics for the endless environments.
Figure 14: Wall-time
We further take a look at the wall-time of various agents trained on Mystery Path (Figure D). It becomes apparent that TrXL and GRU are quite on par concerning wall-time efficiency, whereas TrXL needs fewer samples as shown in section 4.1. On first sight, it seems surprising that GTrXL is faster than TrXL (Table 5) even though GTrXL is more complex. This is due to the more expensive resets of the environment. Better policies have shorter episodes and thus reset more frequently impairing wall-time. So as GTrXL’s and LSTM’s policies are inferior, their wall-time is faster. These results were obtained from the LiDo3 high-performance cluster [^3] https://lido.itmc.tu-dortmund.de/. Each run utilized 32 cores (AMD Epyc 7542), an A100 GPU, and 200GB RAM.
Table 5: Wall-time efficiency metrics for various agents trained on Mytsery Path. Agents are arranged from left to right based on increasing mean training duration, measured in hours. Other values are represented in seconds and comprise the time needed for a single update (i.e. one complete iteration of running PPO). The second row indicates if the neural network was compiled before training.
LSTM GRU GTrXL B0 TrXL
Compiled No No Yes Yes
Mean 5.09 5.19 6.68 7.39
Std 0.22 0.28 0.31 0.24
Min 4.2 4.0 5.8 6.4
Max 7.6 7.0 62.2 11.5
Median 5.1 5.2 6.6 7.4
IQM 5.09 5.18 6.64 7.39
Mean Duration 35.35 36.04 46.4 51.32
Appendix E Agent Behaviors